OpenStudy (anonymous):

find the exact value of the given trigonometric function tan13π/4

OpenStudy (tkhunny):

Reference Angle. What is the reference angle for 13pi/4?

OpenStudy (campbell_st):

ok... so 2pi is 1 revolution \[\frac{13\pi}{4} = \frac{8\pi}{4} + \frac{5\pi}{4} \] and \[\frac{8\pi}{4} = 1 revolution \] so you need to evaluate \[\tan (\frac{5\pi}{4})\] so the angle is in the 3rd quadrant where tan is positive \[\tan(\pi + \frac{\pi}{4}) = \tan(\frac{\pi}{4})\] just find the exact value of \[\tan(\frac{\pi}{4})\]

OpenStudy (anonymous):

find the exact value of the given trigonometric function -cos (π/4)+16π)